OpenStudy (anonymous):

please solve the provided inequality, and make certain to express your solution in: 1. inequality form 2. graphical form, and 3. interval form . 4-x is less than or equal to 8x+2x

OpenStudy (anonymous):

\[4 - x \le 8x + 2x \] \[4 - x \le 10x\] \[4 \le 9x\] \[4/9 \le x\] X is greater than or equal to 4/9. In interval form: \[[4/9, \infty)\]

OpenStudy (anonymous):

so when it asks me to graph what do i graph, idk if he wants us to use the x and y column?

OpenStudy (anonymous):

and how come i got X>=4/11? lol

OpenStudy (anonymous):

yeah, 4/11 is right. I think you just need to graph on a number line.

OpenStudy (anonymous):

ohh okay! thank you